OpenStudy (anonymous):

Solve: x3 – 2x2 – 48x = 0

OpenStudy (anonymous):

Factor out an X. x^2-2x-48.

OpenStudy (anonymous):

x(x2 - 2x - 48) = 0 x(x-8)(x+6)=0 x = 0,8, -6

OpenStudy (anonymous):

Teamwork.

OpenStudy (anonymous):

x(x2 -2x-48) =o consider x(x2 -8x +6x - 48)=0 x=0 but (x2 -8x) +(6x-48)=0 x(x -8) + 6(x-8)=0 (x+6) (x-8)=0 x+6=0.................i x-8=0.................ii x=-6 and x=8 therfore x=0,-6 and 8